### v2

# FOL-Traces: Verified First-Order Logic Reasoning Traces at Scale

Abstract
Reasoning in language models is difficult to evaluate: natural-language traces are unverifiable, symbolic datasets too small, and most benchmarks conflate heuristics with inference. We present FOL-Traces, the first large-scale dataset of programmatically verified reasoning traces, enabling rigorous evaluation of structured logical inference.111We plan to release the full dataset, along with the data generation and analysis code used in this paper. We also propose two challenging and comprehensive diagnostic tasks—masked operation prediction and step completion—that directly probe syntactic awareness and process fidelity. FOL-Traces serves as a scalable testbed for rigorously studying how models perform structured logical inference. Systematic experiments with 5 reasoning LLMs show that the dataset remains challenging: models only reach around 45.7% accuracy on masked operation prediction and around 27% on two-step completion.
FOL-Traces:
Verified
First-Order Logic Reasoning Traces at Scale
Isabelle Lee USC lee.isabelle.g@gmail.com Sarah Liaw Harvard University Dani Yogatama USC
1 Introduction
To establish whether large language models (LLMs) produce correct reasoning, they must be evaluated in domains that provide both structural clarity and verifiable inference. Existing approaches fall short of this requirement. Natural-language chains of thought, while human-interpretable, are poorly defined (bender-koller-2020-climbing), unverifiable, and often unfaithful to the model’s underlying computation tyen2024llmsreasoningerrorscorrect; min2022rethinkingroledemonstrationsmakes. Symbolic datasets such as math problem sets, algebraic puzzles, and theorem-proving corpora provide formal supervision, but they are typically too small or under-annotated for large-scale, systematic reasoning analysis. Widely used benchmarks further conflate heuristic shortcuts with inference, making it difficult to determine whether success reflects systematic reasoning or pattern exploitation.
Recent advances in reasoning-augmented language models—such as OpenAI’s O1 (openai2024o1), DeepSeek’s R1 (deepseekai2025deepseekr1incentivizingreasoningcapability), and the Qwen reasoning model (bai2023qwentechnicalreport)—have intensified interest in structured problem-solving within LLMs. These models have shown strong performance on complex tasks, including AIME mathematical challenges (vals2025aime) and coding benchmarks (chen2021codex; jimenez2024swebench), often leveraging large-scale reinforcement learning and external verifiers. A key strategy is to train on extended sequences that incorporate chains-of-thought (CoT) or reasoning traces (wei2023chainofthoughtpromptingelicitsreasoning). In addition to recent advancements in post-training, emerging research explores how reasoning mechanisms might be integrated into pretraining itself (geiping2025scalingtesttimecomputelatent; tack2025llmpretrainingcontinuousconcepts; ruan2025reasoninglearnlatentthoughts; dong2025reinforcementpretraining; hatamizadeh2025rlpreinforcementpretrainingobjective). Rather than treating reasoning as a byproduct of scale, these approaches suggest it can instead be structured through targeted development. A deeper understanding of how reasoning arises during training could enable us to induce it more deliberately.
1.1 Background on interpretability and Chain-of-Thought (CoT) reasoning
To better understand reasoning in LLMs, a plethora of recent work have turned to mechanistic explanations. lindsey2025biology analyze model circuitry to trace internal reasoning patterns, though their findings are limited to carefully controlled settings. elhage2022toymodelssuperposition, and wang2022interpretabilitywildcircuitindirect identify components like induction heads and algorithmic circuits, offering insights into how models support multi-step reasoning.
However, other recent findings suggest that CoT traces may align poorly with actual inference dynamics. While CoT prompting has improved performance on complex reasoning tasks (wei2023chainofthoughtpromptingelicitsreasoning; kojima2023largelanguagemodelszeroshot), several studies demonstrate that its explanations can be unfaithful to the model’s true computational process (turpin2023languagemodelsdontsay; Wei_Jie_2024). Techniques such as self-consistency sampling (wang2023selfconsistencyimproveschainthought), contrastive CoT (chia2023contrastivechainofthoughtprompting), and faithful reasoning frameworks (lyu2023faithfulchainofthoughtreasoning) attempt to bridge this gap. Nonetheless, the central challenge remains: how to reliably relate surface-level reasoning traces to the internal mechanisms that actually drive model behavior.
These challenges expose a key gap in reasoning datasets. Natural language CoT resources are large but unverifiable due to linguistic ambiguity. Formal corpora such as theorem-proving datasets (e.g., Lean (Yang2023LeanDojoTP), Metamath (yu2024metamathbootstrapmathematicalquestions)) offer symbolic guarantees yet lack detailed annotations or remain too small for modern pretraining. Mathematical datasets like MATH (hendrycksmath2021) and GSM8K (cobbe2021gsm8k) provide intermediate scale but verify only final answers, not reasoning steps. No existing dataset combines programmatic step-level verification with pretraining-scale size. FOL-Traces fills this gap, as shown in Table 1.
| Dataset | Scale | Verified Steps | Domain |
| MATH | 2.93K | ✗ (answers only) | Math |
| GSM8K | 8K | ✗ (answers only) | Math |
| Lean | <100K | ✓ | Formal proofs |
| FOL-TRACES | 7.4M | ✓ | First-order logic |
1.2 Formal Reasoning Presents an Opportunity
Formal reasoning provides a clear lens for examining how language models reason. Unlike natural language reasoning—which is often ambiguous and hard to verify—formal logic offers interpretable, verifiable structure, yet purely symbolic tasks like arithmetic can be too limited and detached from language. First-order logic (FOL) strikes a useful balance: it mirrors natural-language structure while remaining algorithmically grounded and semi-decidable (boolos2007computability)—valid formulas can be confirmed by an algorithm, though invalid ones cannot always be disproved. Leveraging this property, we introduce FOL-Traces, the first large-scale dataset with programmatically verified reasoning traces (7.4M examples, 3.5B tokens). Unlike unverifiable natural language CoT or theorem-proving corpora—which are too small for pretraining—FOL-Traces uniquely combines symbolic guarantees with scale, enabling systematic analysis of how models acquire, represent, and generate logical reasoning with step-by-step verification.
Building on this motivation, our contributions are threefold:
1.
We introduce FOL-Traces, the first large-scale dataset with programmatically verified reasoning traces (7.4M examples, 3.5B tokens)—addressing the critical gap between unverifiable natural language CoT and small-scale symbolic datasets.
2.
We propose two diagnostic tasks—masked operation prediction and step completion—that probe models’ syntactic awareness and process-level reasoning fidelity.
3.
We benchmark both pretrained and reasoning-oriented LLMs, revealing systematic differences in how models acquire and internalize logical structure.
2 FOL-Traces dataset
Most reasoning datasets face a trade-off: natural language CoT is unverifiable, while symbolic logic corpora are too small for large-scale training. FOL-Traces combines programmatic verification with pretraining-scale size, enabling systematic study of logical reasoning. The generation pipeline shown in Figure 1 (1) generates symbolic FOL formulas and (2) instantiates them with natural language predicates and variables via LLMs. We leverage existing benchmarks such as LogicBench parmar2024logicbenchsystematicevaluationlogical and FOLIO han2024folionaturallanguagereasoning as sources of human-annotated logical forms and prompts. In addition to the collected inference rules from existing datasets, we programmatically generate 1.5 million unique rules and corresponding step-by-step simplification traces. Then, we employ LLMs to instantiate them, resulting in a diverse and extensive collection of FOL expressions. We describe the programmatic unique rule generation process in §2.2.1. Then, we detail the various properties and the complexity dynamics of the rule sets and resulting datasets in §3.
2.1 Dataset Structure and Annotations
The dataset contains pretraining-scale training splits, as well as comprehensively labeled validation and test splits. Each example is grounded in an underlying first-order symbolic expression annotated with a unique rule ID. Subsequently, each rule is annotated in detail with the following fields.
Rule (rule). This field contains the first-order logic expression template to be instantiated. Each rule consists of at least one first-order logic formula. For curated rules, we use basic inference rules from first-order logic. For randomly generated rules, we programmatically generate a formula and then simplify it step by step to create a chain of first-order logic expressions that together form a rule.
Expressions (exprs). For randomly generated rules, each formula is simplified using basic first-order logic inference rules. Each simplification step is recorded as an expression and added to the list in this field.
Complexity by step (complexity_by_step). The circuit complexity corresponding to each expression in exprs. Here, we define the circuit complexity of a Boolean or first-order logical formula with
: the set of atomic propositions in (i.e., statements with Boolean values True/False),
: the set of immediate subformulas of , i.e., those that occur exactly one level below in its syntactic tree. For example, if , then .
Each logical connective (e.g., ) corresponds to a logic gate in a Boolean circuit.
Then, we define the circuit complexity as the circuit size needed to compute . Formally, we define as:
where is a logical operator (e.g., ), and . Generally, the circuit complexities decrease for subsequent expressions in exprs, as they simplify.
Elimination complexity(elimination_complexity). Each expression in exprs is a simplification of the one before it. The elimination complexity at index is the program complexity of simplifying exprs[] into exprs[].
Program complexity (program_complexity). This represents the total program complexity involved in randomly generating a first-order logic expression and subsequently, recursively simplifying it. In some cases, the expression is fully simplified. However, if the simplification requires too large a recursive jump, the program may stop prematurely. A detailed description of the full algorithm is provided in §2.2.1.
Original depth (original_depth). The original depth is the depth of the initial expression, i.e., exprs[0].
Finally, Table 2 presents summary statistics for the dataset splits. Compared to other datasets, our reasoning CoT traces are substantially larger by a wide margin. Moreover, all of these existing datasets lack detailed annotations for assessing correctness or for quantifying different types of reasoning failures.
| Dataset | Split | # Examples | # Tokens |
| FOL-Traces | Train | 7.40M | 3.51B |
| Dev | 5K | 3.05M | |
| Test | 8.64K | 5.25M | |
| Curated Train | 8.80M | 786.27M |
2.2 Dataset Generation
We constructed the dataset through a hybrid approach combining human-curated rules, automated symbolic rule generation, and LLM prompting. The detailed breakdown of the inference rules as well as human curated rules are shown in §C, §E, §D, and §F. Our process began with the collection of existing benchmarks for logical reasoning, which provided both direct training material and a basis for generating generalized logical proofs via LLMs.
To expand the diversity of logical forms, we developed templates representing canonical inference rules in first-order logic, including De Morgan’s laws, tautologies, and implications. These templates were instantiated with meaningful predicates using LLMs. For example, the general form
was instantiated as
where and are predicates describing environmental properties. Such instantiations were generated by prompting GPT-4o with the abstract logical form and a set of domain-relevant predicate candidates. Additionally, we used symbolic computation tools such as SymPy to synthesize first-order logic expressions, ensuring both syntactic correctness and logical validity. This multi-stage generation process enabled the creation of a high-quality, diverse dataset designed to evaluate and train models on robust logical reasoning tasks.
2.2.1 Generate rules with SymPy
To generate verifiably correct first-order logic expressions, we utilize SymPy. The generation process follows the algorithm outlined in Algorithm 1, which recursively constructs logical formulas based on a random selection of operators and symbols.
The depth and complexity of the formula are controlled by the parameters provided. Once a formula is generated, a high-level simplification procedure is applied (described in Algorithm 2) to reduce the formula to its simplest form by applying standard logical inference rules such as De Morgan’s laws and eliminating tautologies. The full pseudocode for these processes is provided in §A.1.
Finally, we generate chains of reasoning traces by sequentially linking the simplification steps. This process is akin to Chain-of-Thought (CoT) reasoning, where each intermediate reasoning step builds upon the previous one, forming a structured sequence that leads to the final conclusion. However, unlike traditional reasoning traces, our approach ensures that each step is programmatically verified to be correct, leveraging SymPy’s symbolic computation capabilities. By chaining these simplifications, we capture a stepwise progression of logically sound transformations, which can be useful for tasks that require a detailed, interpretable sequence of logical reasoning.
3 Dataset Statistics
We begin by examining the distribution of rules in our dataset, which were generated to provide a diverse set of templates for constructing complexity-annotated examples used in pretraining. A summary of this distribution is presented in Figure 2.
Figure 2 illustrates the distribution of the circuit complexity , as defined in Equation 1, at step 0. The distribution appears to be a bimodal normal distribution with the mean around 13 and 19. Then, the original circuit complexity in Figure 2 describes . Notably, the distribution appears tri-modal, with a concentration of low-complexity rules around 20, a higher-complexity group near 30, and an overall mean complexity around 35. Then, as the rule simplifies, the total program complexity to generate the rule chains propagates as a trimodal distribution, with mean at 15, 23, and 27.
Next, we examine how rewrite rules evolve throughout the simplification process, as illustrated in Figure 3. The top left panel shows that the number of rewrite steps per rule follows an approximately normal distribution with a long tail, indicating that most rules simplify quickly while a few require many steps. The top right panel depicts elimination complexity at step 0, which is typically low but occasionally high, reflecting rare difficult simplifications. The bottom panel shows that circuit complexity generally increases with rewrite depth, suggesting that longer rewrite sequences correspond to more intricate transformations.
4 Experiments
We conduct two stages of experiments to assess both the stability of our training setup and the reasoning competence it supports. First, we perform pretraining sanity checks to verify that the FOL-Traces corpus enables stable optimization and coherent FOL representations in a small LM. Second, we evaluate reasoning-oriented models on structured diagnostic tasks that probe syntactic awareness and process fidelity, allowing systematic comparison across model families and scales.
4.1 Pretraining Sanity Checks
As a preliminary step, we pretrain Karpathy2022’s implementation of the 125M-parameter GPT-2-small architecture (12 layers, 12 heads per layer) from scratch on the FOL-Traces corpus to validate that the dataset supports stable optimization and coherent FOL representations prior to downstream reasoning evaluations. The model uses an embedding size of 768, a block size of 1024 tokens, and a micro-batch size of 12 with 40 gradient accumulation steps to simulate a larger effective batch. Training employs the AdamW optimizer with a learning rate of , weight decay of 0.1, gradient clipping at 1.0, and no dropout. The learning rate follows a warm-up over 400 iterations and decays to a minimum of across 20,000 iterations. Training was conducted on two NVIDIA RTX A6000 GPUs for approximately 40 hours.
We conduct two complementary studies—one-step True / False evaluation as a basic reasoning task and representation-level probing—to assess how the 125M-parameter GPT-2 variant internalizes FOL.
One-Step True / False Evaluation
As a quick diagnostic, we analyze one-step reasoning with definitive truth values, focusing on the simplest strictly evaluable forms of reasoning. For each formula–interpretation pair , we prompt the model with “” and count a prediction correct when the next token matches the gold True or False. Although the full 50,304 vocabulary is available, the model rapidly collapses its output distribution onto a binary subspace, indicating early acquisition of the truth-evaluation task.
To analyze difficulty, we partition the evaluation set by original complexity into three buckets—low, medium, and high. Figure 11 traces Top-1 accuracy across training tokens, and Table 3 reports the tokens required to reach 70% accuracy on the unseen split as well as the final accuracies. Syntax is learned rapidly: after tokens (5% of training), accuracy improves from 50% to 60%, indicating that the model collapses its output space to a binary decision.
| Bucket | Tokens to 70% | Final (seen) | Final (unseen) |
| 0–21 | 3.2 B | 0.76 | 0.76 |
| 22–32 | 1.1 B | 0.79 | 0.87 |
| 33 | 1.1 B | 0.89 | 0.92 |
Probing Geometric Sensitivity to Syntactic Validity
LLMs are typically evaluated on task accuracy, which overlooks whether their hidden states encode syntactic well-formedness. We recast our True / False corpus into a binary validity benchmark, creating minimal corruptions by deleting one token from fifty True formulas to disrupt global coherence. We compare our pretrained NanoGPT with BERT-base, and compute the cosine similarities between the predicted embeddings and the valid labels. We quantify the Spearman’s bootstrapped over 1,000 samples. NanoGPT’s embeddings showed a much stronger alignment with syntactic validity () than BERT’s (), likely because GPT’s autoregressive training enforces a kind of “syntactic checksum” that quickly detects missing or malformed tokens, whereas BERT’s bidirectional masking spreads error signals across the sequence, weakening its sensitivity to such structural breaks. Overall, this compact FOL corpus and embedding-level RSA reveal that a small LM can encode syntactic structure geometrically when trained with FOL-Traces.
| Model | Representation | 95% CI | |
| NanoGPT (20 k) | final BOS stream | 0.682 | [0.666, 0.695] |
| BERT-base | final [CLS] | 0.327 | [0.308, 0.345] |
4.2 Structured Reasoning Diagnostic Tasks
We introduce and perform two diagnostic tasks—masked operation prediction and step completion—that directly probe syntactic awareness and process fidelity. These tasks are designed to isolate core reasoning behaviors: the masked operation prediction task targets atomic-level agreement in first-order logic syntax, while the step completion task evaluates the model’s ability to extend reasoning chains consistent with first-order logic. Both tasks are automatically evaluable via programmatic verifiers, enabling scalable and objective assessment. Experiments are conducted using five reasoning models from two families, Phi-3 (abdin2024phi3technicalreporthighly) and Qwen-2.5 (bai2023qwentechnicalreport), covering a range of model sizes.
4.2.1 Task 1: Masked Operation Prediction
For a targeted evaluation of distinct aspects of compositional reasoning, we isolate specific dimensions of the model’s compositional understanding through targeted input masking. As illustrated in Figure 5, Task 1 employs three masking schemes to separate reasoning components: component masks (blue) obscure subexpressions or entities, operator masks (green) conceal logical or relational connectors, and predicate masks (red) hide function or action tokens. This setup enables controlled analysis of the model’s internal representations, revealing how each subtask contributes to the overall reasoning process. The masked operation prediction tasks probe distinct facets of compositional reasoning: component accuracy measures the model’s ability to correctly identify and process complete subexpressions within logical expressions; operator accuracy evaluates whether the model correctly applies logical or relational operators between components; and predicate accuracy assesses correctness in identifying function names. Together, these metrics provide a holistic view of the model’s systematic reasoning capabilities.
Across all evaluated models, performance varies considerably by both model family and size. Within Phi-3 models, accuracy remains relatively low across all metrics, with phi-3-medium-128k performing particularly poorly and phi-3.5-mini showing only modest gains. In contrast, the Qwen-2.5 series demonstrates a clear scaling trend: larger models consistently outperform smaller ones, culminating in qwen2.5-32b, which achieves the highest scores in every category—75% component accuracy, 73% operator accuracy, and 80% predicate accuracy. Despite these improvements, even the strongest models show limited robustness and generalization across compositional and operational reasoning tasks. Notably, models explicitly trained for reasoning, such as phi-3.5-mini, still underperform substantially compared to larger general-purpose models. This suggests that current reasoning-oriented training approaches alone may be insufficient for achieving strong systematic reasoning performance, highlighting the need for more integrated reasoning and compositional generalization strategies.
| Model | Comp. Acc | Oper. Acc | Pred. Acc | Overall |
| phi-3.5-mini | 26.00% | 32.00% | 19.00% | 25.67% |
| phi-3-medium-128k | 1.00% | 22.00% | 17.00% | 13.33% |
| qwen2.5-7b | 64.00% | 25.00% | 62.00% | 50.33% |
| qwen2.5-14b | 72.00% | 42.00% | 75.00% | 63.00% |
| qwen2.5-32b | 75.00% | 73.00% | 80.00% | 76.00% |
| Overall | 47.60% | 38.80% | 50.60% | 45.67% |
4.2.2 Task 2: Step Completion
For the step completion task, we ask the models to complete the last couple steps given a FOL chain. Using SymPy as an automatic verifier, we programmatically check whether the generated reasoning steps are logically equivalent to the gold steps. Table 6 summarizes the results, showing that the Qwen2.5 models outperform Phi-3. However, even the best-performing Qwen2.5-32B model struggles to surpass 50% accuracy, indicating that this step completion task remains highly challenging for large reasoning models.
| Model | 1-Step | 2-Step | 2-Step Chain |
| phi-3.5-mini | 0.10 | 0.00 | 0.00 |
| phi-3-medium-128k | 0.10 | 0.08 | 0.20 |
| qwen2.5-7b | 0.54 | 0.22 | 0.44 |
| qwen2.5-14b | 0.60 | 0.26 | 0.48 |
| qwen2.5-32b | 0.76 | 0.40 | 0.48 |
| Overall | 0.43 | 0.16 | 0.27 |
A more detailed break down for 2-steps completion is shown in Figure 6. We see that all the models perform on the last step prediction the better than the two step prediction.
We further stratify the results based on the original complexities of the underlying symbolic FOL rules, as shown in Figure 7. For both the last 1- and 2-step completion tasks, LLMs achieve higher accuracy on low-complexity rules (yellow), while their performance declines markedly as rule complexity increases. As illustrated in Figure 7(a), several reasoning LLMs are able to correctly generate low-complexity rules; however, their accuracy drops substantially for medium (green)- and high (red)-complexity examples. Similarly, Figure 7(b) shows that recent reasoning LLMs perform poorly even on relatively simple cases—with the best-performing model, Qwen2.5-32B, attaining only around 50% accuracy for low- and medium-complexity examples. Most models fail to generate correct outputs for high-complexity cases (with the exception of Qwen2.5-14B, which reaches approximately 20% accuracy), underscoring that this task—and derivative -step completion tasks—would remain highly challenging for current LLMs.
Figure 8 shows the error breakdown by model. Most errors stem from step 1 (blue), step 2 (yellow), or both (red). Additional failures include formatting issues (grey, where the generated FOL expression was malformed and unparseable by SymPy) and chain-only errors, where the model captures general, overarching logical relations of the chain but introduces structural mistakes such as subexpression swaps or misplaced predicates. An example of chain-only errors are shown in Table 14.
Our analyses indicate that reasoning LLMs may not be reasoning as rigorously as expected. They still perform poorly on both tasks 1 and 2, with error patterns revealing fundamental misunderstandings of basic logical syntax.
5 Conclusion
We introduce FOL-Traces, the first large-scale dataset of programmatically verified reasoning traces. In contrast to unverifiable natural language CoT corpora and limited theorem-proving datasets, FOL-Traces provides 7.4M verified examples with fine-grained complexity annotations, addressing a critical gap in reasoning evaluation and training. To assess its utility, we conduct pretraining sanity checks confirming stable optimization and coherent logical representations, followed by two structured reasoning diagnostics: masked operation prediction and step completion. Qwen2.5-32B attains 75% accuracy on component prediction, while reasoning-tuned models such as Phi-3.5-mini perform markedly worse; on two-step completions, all models remain below 50%. These results expose persistent limitations in compositional and process-level reasoning, establishing FOL-Traces as a scalable and verifiable benchmark for advancing genuine reasoning competence beyond surface-level pattern matching.
6 Limitations and Future Work
While FOL-Traces offers a transparent and verifiable foundation for studying structured reasoning in language models, several limitations remain. The dataset focuses on first-order logic, which, though expressive and formally grounded, captures only part of the reasoning diversity found in naturalistic contexts. Its symbolic generation process, while rigorously validated, diverges from the distributional properties of typical pretraining corpora, potentially limiting transferability to open-domain reasoning. Moreover, our evaluations target diagnostic reasoning tasks rather than full downstream applications, leaving generalization and scaling behavior for future study. Despite these constraints, FOL-Traces provides a controlled platform for probing symbolic competence and advancing more interpretable and transparent reasoning systems.
Appendix A Appendix
A.1 Full Simplification Algorithm
Appendix B Additional Dataset Distribution Information
| Bucket | #rules | Start | End | |
| 2 | 82,490 | 15.59 | 5.76 | +9.83 |
| 4 | 217,298 | 16.06 | 7.23 | +8.83 |
| 8 | 225,177 | 17.42 | 9.61 | +7.82 |
| 12 | 87,316 | 19.91 | 11.75 | +8.16 |
| 15 | 25,925 | 22.44 | 13.65 | +8.79 |
| 20 | 1,451 | 25.42 | 14.68 | +10.74 |
| 25 | 15 | 28.00 | 13.80 | +14.20 |
Appendix C First Order Logic (FOL) Categories and Explanations
| FOL Inference Rule | Symbolic Expression | Explanation | |||
| Bidirectional Dilemma (BD) |
| Constructive Dilemma (CD) |
| Destructive Dilemma (DD) |
| Disjunctive Syllogism (DS) |
| Hypothetical Syllogism (HS) |
| Modus Ponens (MP) |
| Modus Tollens (MT) |
| Universal Instantiation (UI) |
| Existential Generalization (EG) |
| FOL proofs & general statements | – | – |
| FOL Properties | Symbolic Expression | |
| Distributive (Dist) |
| Association (AS) |
| Tautology (TT) |
| Transposition (TS) |
| Importation (IM) |
| Exportation (EX) |
| Double Negation (DN) |
| De Morgan’s Law (DM) |
| Negation of XOR (NX) |
| Negation of XNOR (NN) |
Appendix D Elimination Rules
| Symbolic Expression | ||
| E0 |
| E1 |
| E2 |
| E3 |
| E4 |
| E5 |
| E6 |
| E7 |
| E8 |
| E9 |
| E10 |
| E11 |
| E12 |
| E13 |
| E14 |
| E15 |
| E16 |
| E17 |
| E18 |
| E19 |
| E20 |
| E21 |
| E22 |
| E23 |
| E24 |
| E25 |
| E26 |
| E27 |
| E28 |
| E29 |
| E30 |
| E31 |
| E32 |
| E33 |
Appendix E Complex FOL Expressions
| C1 |
| C2 |
| C3 |
| C4 |
| C5 |
| C6 |
| C7 |
| C8 |
| C9 |
| C10 |
| C11 |
| C12 |
| C13 |
| C14 |
| C15 |
| C17 |
| C18 |
| C20 |
| C21 |
| C22 |
| C23 |
Appendix F Training Data
| # Examples | # Tokens | # Examples | # Tokens | ||
| BD | 102.43K | 778.57K | DM | 740.92K | 34.11M |
| CD | 856.78K | 71.70M | Dist | 268.76K | 17.97M |
| DD | 806.15K | 71.70M | XOR∗ | 17.46K | 793.75K |
| DS | 321.31K | 12.79M | XNOR∗ | 15.42K | 668.07K |
| HS | 429.94K | 23.20M | XOR-XNOR∗ | 14.49K | 577.84K |
| MP | 237.80K | 7.49M | |||
| MT | 285.19K | 11.10M | |||
| UI | 123.00K | 4.03M | |||
| EG | 9.10K | 263.0K | |||
| General/fol proof | 2.27M | 286.12M | |||
| C2 | 94.36K | 5.26M | E0 | 5.23K | 83.92K |
| C3 | 108.56K | 20.21M | E1 | 5.34K | 141.32K |
| C4 | 102.73K | 8.03M | E2 | 5.28K | 147.74K |
| C5 | 107.64K | 16.67M | E3 | 5.46K | 88.60K |
| C7 | 109.03K | 17.57M | E4 | 5.18K | 82.91K |
| C8 | 97.90K | 12.38M | E5 | 5.31K | 146.08K |
| C10 | 102.43K | 3.84M | E6 | 5.35K | 148.27K |
| C11 | 86.92K | 3.55M | E7 | 5.28K | 84.61K |
| C13 | 93.22K | 5.17M | E8 | 5.40K | 188.11K |
| C14 | 109.20K | 10.03M | E9 | 5.37K | 194.38K |
| C17 | 108.53K | 10.04M | E10 | 5.23K | 140.78K |
| C20 | 109.10K | 10.58M | E11 | 5.22K | 139.83K |
| C21 | 109.37K | 11.21M | E12 | 5.16K | 134.20K |
| C23 | 109.45K | 14.94M | E13 | 5.18K | 135.07K |
| E14 | 5.60K | 135.07K | |||
| E15 | 5.58K | 720.92K | |||
| E16 | 5.56K | 441.38K | |||
| E17 | 5.58K | 569.09K | |||
| E18 | 5.61K | 256.89K | |||
| E19 | 5.37K | 291.75K | |||
| E20 | 5.21K | 286.32K | |||
| E21 | 5.37K | 352.57K | |||
| E22 | 5.61K | 734.95K | |||
| E23 | 5.59K | 470.14K | |||
| E24 | 5.63K | 586.41K | |||
| E25 | 5.58K | 703.58K | |||
| E26 | 5.57K | 714.06K | |||
| E27 | 5.58K | 501.50K | |||
| E28 | 5.57K | 502.50K | |||
| E29 | 5.59K | 625.27K | |||
| E30 | 5.59K | 720.77K | |||
| E31 | 5.59K | 499.31K | |||
| E32 | 5.59K | 498.49K | |||
| E33 | 5.60K | 648.88K | |||
| Complex Total | 1.45M | 149.48M | Eliminations Total | 185.09K | 12.24M |
| Random | 15.26M | 1.89B | |||
| Total | 24.07M | 2.67B |
*We explicitly included negations of XOR, negations of XNOR and the equivalences between XOR and XNOR.
Appendix G Sanity Check True /False Experiment Plot
Appendix H Task 2 Prompts
| 1-Step Prompt | 2-Step Prompt |
You are given a first-order logic equivalence chain where the last step is missing (replaced by <BLANK>).
[…]
You are given a first-order logic equivalence chain where the last 2 steps are missing (replaced by <BLANK>).
[…]
Appendix I Task 2 Example
| Examples | Expressions and Evaluation Results |
| Step-2 correct, chain wrong |
Answer:
Prediction: (both end with False) [] |
| Different Step-2, chain still correct |
Answer:
Prediction: (both end with False) [] |